OpenStudy (anonymous):

Alice was biking at 8mph. After 3minutes, Bob started after Alice and it took 7minutes to catch up with Alice. How fast was Bob biking?

OpenStudy (anonymous):

Look at it this way, converted to minutes.....\[8 \frac{mi}{60\min}(10 \min)=x \frac{mi}{60\min}(7\min) \implies (10)\frac{8}{7}=\frac{80}{7}=x\]

OpenStudy (anonymous):

I guess if you convert to decimal notation, you will get a little over 11.4 mph.